Colin invested 10,000, part at 14% and part at 17%. If the total interest at the end of the year is 1,490, how much did he invest at 14%?
Answer by checkley77(12569)

.14x+.17(10,000-x)=1,490
.14x+1,700-.17x=1,490
-.03x=1,490-1,700
-.03x=-210
x=210/.03
x=$7,000 invested @ 14%.
10,000-7,000=3,000 invested @ 17% Proof:
.14*7,000+.17*3,000=1,490
980+510=1,490
1,490=1,490